A corporation has an outstanding bond with the following characteristics:

Coupon Rate 6.0%
Interest Payments Semi-Annually
Face Value $1,000
Years to maturity 20
Current market value $854.64

What is the yield to maturity (YTM) for this bond?

**Please show how you got the answer**
Answer #1

Using financial calculator
PMT=6%*1000/2=30
FV=1000
PV=-854.64
N=20*2=40
CPT I/Y=3.70%

Hence, yield to maturity YTM=3.70%*2=7.40%